# Knowledge Enhanced Conditional Imputation for Healthcare Time-series

### v4

@@ I Introduction
Multivariate time series obtained from electronic health records (EHR) play a crucial role in predictive healthcare analytics to derive patient-specific insight [1, 2, 3]. However, EHR time series are characterised by their complexity, and designing successful downstream predictive models requires addressing the abundance of missingness within correlated variables recorded over time. EHR missingness is a direct consequence of clinical and administrative workflows, which inherently lead to variations in the frequency and timing of documenting vital signs, laboratory tests, and other healthcare measures. For example, heart rate is routinely monitored, while data on white blood cell counts are ordered in specific clinical scenarios, such as suspected infection. Furthermore, the structure of EHR data gives rise to numerous correlations between the feature values and their missingness patterns over time. For example, because hypertension is a known cause of kidney disease, not only are high blood pressure measurements expected along with abnormal creatinine levels, but EHR datasets generally show high correlations between the recording patterns of blood pressure and urine creatinine levels [4]. These inherent irregularities render more than 50% of EHR data missing not at random [5], with significant variations in missingness patterns between tasks and datasets [5]. These challenges create a complex landscape for imputation algorithms [6, 7, 8, 9, 10, 11, 12] and highlight the importance of domain knowledge in improving modelling accuracy.
@@
Recent advances in deep learning have shown significant potential in addressing the complexities of EHR missingness. In particular, models based on recurrent neural networks (RNNs), such as GRU-D [13], BRITS [14], and M-RNN [15], have proven effective in capturing the temporal and cross-sectional dependencies embedded within EHR data. Among these, BRITS has shown superior performance in datasets of varying characteristics [16, 17]. BRITS uses bi-directional RNN dynamics and treats missing data as variables of the bidirectional RNN graph, subsequently involving them in the backpropagation process.
@@
Despite advances, none of the existing models, including BRITS, account for domain-imposed disparities in feature-recording patterns, which can have critical implications on imputation accuracy. We show that incorporating a domain-informed strategy to capture data recording patterns is particularly suited for our domain. In addition, the temporal correlations in the EHR data can span both short- and long-term intervals. For example, while monitoring diabetic patients, both short-term glucose fluctuations and long-term HbA1c levels must be accurately modelled, while cardiovascular risk assessment requires careful cross-sectional examination of biomarkers like cholesterol, glucose, and blood pressure. Although RNNs excel at capturing short-term temporal dependencies in a time series, they are less capable of capturing long-range correlations over time [8]. Finally, all current models rely on random masking as a pre-processing step to generate ground truths for evaluation purposes. This involves designating a randomly selected subset of complete values as missing, primarily creating missing completely at random (MCAR) scenarios and drastically oversimplifying the temporal and cross-sectional dependencies in real EHR time series.
@@ II Related Work
The ability of RNNs to capture temporal evolution has made them an attractive choice for time-series imputation. The GRU-D model [13] introduced the use of temporal decay to recurrent imputation models, whereby the influence of past recordings on missing data decreases over time, with features imputed closer to default values if their last observation occurred a long time ago. Using this concept, GRU-D estimates missing values through a combination of the last recorded value and the global mean within a recurrent architecture. Although this approach alone has shown impressive performance in certain healthcare datasets, its effectiveness has not generalised in others [14], particularly with respect to accounting for discrepancies in vital signs recording frequencies and laboratory tests.
@@
Several extensions of GRU-D exist. M-RNN [15] and BRITS [14] use bi-directional RNNs to capture temporal dynamics in both forward and backward directions, improving imputation accuracy. Both M-RNN afnd BRITS combine RNNs with a feed-forward component to capture additional correlations across features. However, M-RNN treats imputed values as fixed constants while BRITS treats missing values as variables, updating them through backpropagation. Using this strategy, missing values accumulate gradients in both forward and backward directions with consistency constraints, increasing the accuracy of estimated values. BRITS has demonstrated superior performance across various domains [16, 18].
@@
Transformers have shown strong performance in capturing long-term dependencies through their self-attention mechanism [19] and have been used to model missingness at random in SAITS [20]. However, Transformers face challenges in capturing short-term dynamics in time series data and their effectiveness in handling fine-grained temporal details and complex missing patterns remains debatable [19]. Models like NRTSI [21], though designed to handle these issues, struggle with computational complexity [20].
@@
Some imputation efforts prioritise the ability to quantify one’s confidence in the resulting imputation, introducing stochasticity into their architectures. This includes RNN-based models such as GRU-U [22], variational autoencoders such as V-RIN [23], generative adversarial networks such as $E^{2}$ [24] and diffusion models as in CSDI [25]. Apart from CSDI, all these approaches require additional stochastic modules added to the imputing network, resulting in unstable training due to increased coupling, and ultimately leading to less accurate imputation in comparative studies [26]. Although CSDI [25] successfully exploits the correlations between observed values to measure stochasticity, it suffers from data leakage as indicated in its repository.
@@
CSAI aims to build on the current state of neural network imputation by addressing unsolved challenges. By encoding transformer-based self-attention to a bi-directional RNN architecture, CSAI captures long-term dependencies which may be missed by BRITS and similar RNN-based models [14]. Furthermore, introducing the domain-informed temporal decay function refines the imputation process beyond standard decay techniques [13]. Finally, integrating CSAI with a non-uniform masking strategy specifically targets the complex temporal and cross-sectional relationships in EHRs.
@@ III-A Incomplete Multivariate Time-series Representation
[…]
@@ IV-A EHR-Tailored BRITS Adaptations
[…]
@@ IV-B Non-Uniform Masking Strategy
  Input: Dataset $\boldsymbol{X}$ with $D$ features, masking rate $U$, weighting parameter $I$
@@ V-A Datasets
Dataset Size Avg Baseline Missingness Feature Correlation (Static; Categorical) Variables
@@
    eICU [^1] https://physionet.org/content/eicu-crd/2.0/ [27] is a multicentre database with anonymised patient records from intensive care units (ICU) obtained from more than 200 hospitals. eICU is publicly available after registration. We followed the only benchmark extract available for eICU [28].
@@
    MIMIC-III [^2] https://physionet.org/content/mimiciii/1.4/ [29] is an extensive, single-hospital database of over 40,000 critical care patients. We used MIMIC-III to increase the heterogeneity of our datasets with different features types and data dimensionality and followed a well-cited benchmark [30] to produce a dataset comprising 59 numerical variables for 21,128 samples.
@@
    PhysioNet Challenge 2012 dataset is a public medical benchmarking dataset provided by [31], containing records of 4000 48-hour hospital stays, allowing unbiased evaluations of different model performances. This benchmark comprises 35 numerical variables for 3,997 samples.
@@ V-B Experimental Design
    Experiment I: compares CSAI with the best-performing RNN models using the three benchmark datasets and three masking rations (5%, 10% and 20%). Here, CSAI is compared with our backbone bidirectional RNN model BRITS [14], the Gated Recurrent Unit with Decay (GRU-D) [13], the Variational Recurrent Imputation Network (V-RIN) [23], and the Multi-directional Recurrent Neural Network (M-RNN) model [15]. In this experiment, two tasks were undertaken to ensure a fair comparison: 1) because the original BRITS implementation uses LSTM cells while all other baseline models use GRU cells, we implemented a GRU-based BRITS model to ensure a balanced evaluation across similar architectures, including its output along with the original BRITS (BRITS-GRU in the results), 2) we employed our nonuniform masking algorithm as a pre-processing step for all models to avoid any bias arising from the masking step.
@@
    Experiment II: is a large-scale experimental evaluation with different flavours of neural imputation models using the Physionet dataset with a 10% masking ratio. This experiment was conducted using the PyPOTS library [^3] https://pypots.com/about/ [32], an open-source Python toolkit for benchmarking and analysing incomplete multivariate time series. PyPOTS provides access to 29 imputation algorithms, including CSAI and a number of benchmark datasets, including Physionet. In this experiment, CSAI is compared against a broad range of neural network imputation models, such as Transformers, CNNs, GNNs, and diffusion models, as well as RNNs and standard statistical imputation methods. As non-uniform masking is not yet integrated into PyPOTS, we rely on the available masking setup, aiming to evaluate the contribution of CSAI’s temporal decay and attention mechanism on the resulting performance.
    The reader should note that although a comparison with generative models through $E^{2}$ [33] would have provided valuable insight, particularly given its lack of its quantitative comparison with BRITS, the publicly available implementation of $E^{2}$ is incompatible with our current setup and is also not part of PyPOTS.
@@ V-C Experimental Setup
Dataset Pre-processing and Missingness Patterns Simulation: Each dataset was normalised to zero mean and unit variance. For each of the four datasets, we created three versions by further masking 5%, 10% and 20% cells in addition to the missingness already present in the original datasets. These masked cells have known ground truths and will form the basis of our comparison of imputation performance. To more effectively simulate EHR missingness scenarios, our first set of experiments use our non-uniform masking as a pre-processing step to enable capturing patterns representative of common missingness data mechanisms in EHRs to the masked cells. In our wider experiments using PyPOTs, we used the package’s PyGrinder toolbox  [^4] https://pypots.com/ecosystem/#PyGrinder, which implements structured missingness patterns that reflect the dynamic nature of time series data [34, 35, 36].
@@
Evaluation Metrics: To assess imputation performance, we use mean absolute error (MAE) [14] defined as:

### v5

@@ I Introduction
Multivariate time series obtained from electronic health records (EHR) play a crucial role in predictive healthcare analytics to derive patient-specific insight [healthcare, kdop, farah]. However, EHR time series are characterised by their complexity, and designing successful downstream predictive models requires addressing the abundance of missingness within correlated variables recorded over time. EHR missingness is a direct consequence of clinical and administrative workflows, which inherently lead to variations in the frequency and timing of documenting vital signs, laboratory tests, and other healthcare measures. For example, heart rate is routinely monitored, while data on white blood cell counts are ordered in specific clinical scenarios, such as suspected infection. Furthermore, the structure of EHR data gives rise to numerous correlations between the feature values and their missingness patterns over time. For example, because hypertension is a known cause of kidney disease, not only are high blood pressure measurements expected along with abnormal creatinine levels, but EHR datasets generally show high correlations between the recording patterns of blood pressure and urine creatinine levels [correlation]. These inherent irregularities render more than 50% of EHR data missing not at random [wells2013strategies], with significant variations in missingness patterns between tasks and datasets [wells2013strategies]. These challenges create a complex landscape for imputation algorithms [oursurvey, yadav2018mining, jensen2012mining, survey, hu2017strategies, tsibench, unveiling] and highlight the importance of domain knowledge in improving modelling accuracy.
@@
Recent advances in deep learning have shown significant potential in addressing the complexities of EHR missingness. In particular, models based on recurrent neural networks (RNNs), such as GRU-D [grud], BRITS [cao2018brits], and M-RNN [yoon2017multi], have proven effective in capturing the temporal and cross-sectional dependencies embedded within EHR data. Among these, BRITS has shown superior performance in datasets of varying characteristics [survey1, survey2]. BRITS uses bi-directional RNN dynamics and treats missing data as variables of the bidirectional RNN graph, subsequently involving them in the backpropagation process.
@@
Despite advances, none of the existing models, including BRITS, account for domain-imposed disparities in feature-recording patterns, which can have critical implications on imputation accuracy. We show that incorporating a domain-informed strategy to capture data recording patterns is particularly suited for our domain. In addition, the temporal correlations in the EHR data can span both short- and long-term intervals. For example, while monitoring diabetic patients, both short-term glucose fluctuations and long-term HbA1c levels must be accurately modelled, while cardiovascular risk assessment requires careful cross-sectional examination of biomarkers like cholesterol, glucose, and blood pressure. Although RNNs excel at capturing short-term temporal dependencies in a time series, they are less capable of capturing long-range correlations over time [jensen2012mining]. Finally, all current models rely on random masking as a pre-processing step to generate ground truths for evaluation purposes. This involves designating a randomly selected subset of complete values as missing, primarily creating missing completely at random (MCAR) scenarios and drastically oversimplifying the temporal and cross-sectional dependencies in real EHR time series.
@@ II Related Work
The ability of RNNs to capture temporal evolution has made them an attractive choice for time-series imputation. The GRU-D model [grud] introduced the use of temporal decay to recurrent imputation models, whereby the influence of past recordings on missing data decreases over time, with features imputed closer to default values if their last observation occurred a long time ago. Using this concept, GRU-D estimates missing values through a combination of the last recorded value and the global mean within a recurrent architecture. Although this approach alone has shown impressive performance in certain healthcare datasets, its effectiveness has not generalised in others [cao2018brits], particularly with respect to accounting for discrepancies in vital signs recording frequencies and laboratory tests.
@@
Several extensions of GRU-D exist. M-RNN [yoon2017multi] and BRITS [cao2018brits] use bi-directional RNNs to capture temporal dynamics in both forward and backward directions, improving imputation accuracy. Both M-RNN afnd BRITS combine RNNs with a feed-forward component to capture additional correlations across features. However, M-RNN treats imputed values as fixed constants while BRITS treats missing values as variables, updating them through backpropagation. Using this strategy, missing values accumulate gradients in both forward and backward directions with consistency constraints, increasing the accuracy of estimated values. BRITS has demonstrated superior performance across various domains [survey1, britshigh].
@@
Transformers have shown strong performance in capturing long-term dependencies through their self-attention mechanism [transformerstimeseriessurvey] and have been used to model missingness at random in SAITS [du2023saits]. However, Transformers face challenges in capturing short-term dynamics in time series data and their effectiveness in handling fine-grained temporal details and complex missing patterns remains debatable [transformerstimeseriessurvey]. Models like NRTSI [10095054], though designed to handle these issues, struggle with computational complexity [du2023saits].
@@
Some imputation efforts prioritise the ability to quantify one’s confidence in the resulting imputation, introducing stochasticity into their architectures. This includes RNN-based models such as GRU-U [jun2020uncertainty], variational autoencoders such as V-RIN [mulyadi2021uncertainty], generative adversarial networks such as $E^{2}$GAN [luo2018multivariate] and diffusion models as in CSDI [tashiro2021csdi]. Apart from CSDI, all these approaches require additional stochastic modules added to the imputing network, resulting in unstable training due to increased coupling, and ultimately leading to less accurate imputation in comparative studies [mescheder2018training]. Although CSDI [tashiro2021csdi] successfully exploits the correlations between observed values to measure stochasticity, it suffers from data leakage as indicated in its repository.
@@
CSAI aims to build on the current state of neural network imputation by addressing unsolved challenges. By encoding transformer-based self-attention to a bi-directional RNN architecture, CSAI captures long-term dependencies which may be missed by BRITS and similar RNN-based models [cao2018brits]. Furthermore, introducing the domain-informed temporal decay function refines the imputation process beyond standard decay techniques [grud]. Finally, integrating CSAI with a non-uniform masking strategy specifically targets the complex temporal and cross-sectional relationships in EHRs.
@@ III-A Incomplete Multivariate Time-series Representation
Figure 1: An example of multivariate time-series. Observations $\boldsymbol{x_{1-5}}$ in time-stamps $\boldsymbol{s_{1-5}}=0,\,4,\,5,\,7,\,9$. Feature $d_{2}$ was missing during $\boldsymbol{s_{2-4}}$, the last observation took place at $\boldsymbol{s_{1}}$. Hence, $\boldsymbol{\delta_{5}^{2}}=\boldsymbol{t_{5}-t_{1}}=9-0=9$.
@@ IV-A EHR-Tailored BRITS Adaptations
$\boldsymbol{H_{1}}$ $=\text{Conv1D}_{1}(\boldsymbol{C}_{\text{out}}\boldsymbol{W_{1}}+\boldsymbol{b_{1}})$ (19)
$\boldsymbol{h}_{\text{init}}$ $=\text{Conv1D}_{2}(\boldsymbol{H_{1}}\boldsymbol{W_{2}}+\boldsymbol{b_{2}})$ (20)
@@ IV-B Non-Uniform Masking Strategy
Input: Dataset $\boldsymbol{X}$ with $D$ features, masking rate $U$, weighting parameter II
@@ V-A Datasets
Dataset Size
Avg Baseline
Missingness
Feature
Correlation
(Static; Categorical)
Variables
@@
    eICU [^1] https://physionet.org/content/eicu-crd/2.0/ [pollard2018eicu] is a multicentre database with anonymised patient records from intensive care units (ICU) obtained from more than 200 hospitals. eICU is publicly available after registration. We followed the only benchmark extract available for eICU [eicubenchmark].
@@
    MIMIC-III [^2] https://physionet.org/content/mimiciii/1.4/ [johnson2016mimic] is an extensive, single-hospital database of over 40,000 critical care patients. We used MIMIC-III to increase the heterogeneity of our datasets with different features types and data dimensionality and followed a well-cited benchmark [harutyunyan2019multitask] to produce a dataset comprising 59 numerical variables for 21,128 samples.
@@
    PhysioNet Challenge 2012 dataset is a public medical benchmarking dataset provided by [physionetchallenge], containing records of 4000 48-hour hospital stays, allowing unbiased evaluations of different model performances. This benchmark comprises 35 numerical variables for 3,997 samples.
@@ V-B Experimental Design
    Experiment I: compares CSAI with the best-performing RNN models using the three benchmark datasets and three masking rations (5%, 10% and 20%). Here, CSAI is compared with our backbone bidirectional RNN model BRITS [cao2018brits], the Gated Recurrent Unit with Decay (GRU-D) [grud], the Variational Recurrent Imputation Network (V-RIN) [mulyadi2021uncertainty], and the Multi-directional Recurrent Neural Network (M-RNN) model [yoon2017multi]. In this experiment, two tasks were undertaken to ensure a fair comparison: 1) because the original BRITS implementation uses LSTM cells while all other baseline models use GRU cells, we implemented a GRU-based BRITS model to ensure a balanced evaluation across similar architectures, including its output along with the original BRITS (BRITS-GRU in the results), 2) we employed our nonuniform masking algorithm as a pre-processing step for all models to avoid any bias arising from the masking step.
@@
    Experiment II: is a large-scale experimental evaluation with different flavours of neural imputation models using the Physionet dataset with a 10% masking ratio. This experiment was conducted using the PyPOTS library [^3] https://pypots.com/about/ [pypots], an open-source Python toolkit for benchmarking and analysing incomplete multivariate time series. PyPOTS provides access to 29 imputation algorithms, including CSAI and a number of benchmark datasets, including Physionet. In this experiment, CSAI is compared against a broad range of neural network imputation models, such as Transformers, CNNs, GNNs, and diffusion models, as well as RNNs and standard statistical imputation methods. As non-uniform masking is not yet integrated into PyPOTS, we rely on the available masking setup, aiming to evaluate the contribution of CSAI’s temporal decay and attention mechanism on the resulting performance.
    The reader should note that although a comparison with generative models through $E^{2}$GAN [luo2019e2gan] would have provided valuable insight, particularly given its lack of its quantitative comparison with BRITS, the publicly available implementation of $E^{2}$GAN is incompatible with our current setup and is also not part of PyPOTS.
@@ V-C Experimental Setup
Dataset Pre-processing and Missingness Patterns Simulation: Each dataset was normalised to zero mean and unit variance. For each of the four datasets, we created three versions by further masking 5%, 10% and 20% cells in addition to the missingness already present in the original datasets. These masked cells have known ground truths and will form the basis of our comparison of imputation performance. To more effectively simulate EHR missingness scenarios, our first set of experiments use our non-uniform masking as a pre-processing step to enable capturing patterns representative of common missingness data mechanisms in EHRs to the masked cells. In our wider experiments using PyPOTs, we used the package’s PyGrinder toolbox  [^4] https://pypots.com/ecosystem/#PyGrinder, which implements structured missingness patterns that reflect the dynamic nature of time series data [mitra2023learning, cini2022grin, khayati2020mind].
@@
Evaluation Metrics: To assess imputation performance, we use mean absolute error (MAE) [cao2018brits] defined as: